OpenStudy (anonymous):

simplify the radical expression. square root of 39 (square root of 6 + 7)

OpenStudy (anonymous):

\[\sqrt{39}(\sqrt{6} +7)\]

OpenStudy (anonymous):

This is the same as the last one, can you try to use the method I used in the last one to solve this one?

OpenStudy (campbell_st):

\[\sqrt{39}(\sqrt{6}+7) = \sqrt{39\times6} + 7\times \sqrt{39}\] simplifying its \[\sqrt{234} + 7\sqrt{39}\] which can be simplified further \[3\sqrt{26}+7\sqrt{39}\]